# CREST: Cross-modal Resonance through Evidential Deep Learning for Enhanced Zero-Shot Learning

### v2

@@ Abstract.
ccs: Computing methodologies Learning paradigms22footnotetext: Corresponding author.
@@ 1. Introduction
Humans frequently possess the talent to grasp novel concepts relying on prior experience without the need to see them beforehand. For instance, a peacock is commonly known as a bird with a colorful fan-shaped tail; if individuals have previous knowledge of birds and fans, they can quickly identify a peacock. However, unlike humans, widely used and studied supervised deep learning models are typically limited to classifying samples belonging to categories seen during training, lacking the capacity to handle samples from unseen categories during training, thus lacking generality and flexibility. Therefore, to further advance Artificial General Intelligence (AGI) (Bostrom, 2020) and achieve true implementation, Zero-Shot Learning (ZSL) was introduced to identify new classes by leveraging inherent semantic relationships during learning (Larochelle et al., 2008; Palatucci et al., 2009; Lampert et al., 2009, 2013; Fu et al., 2017). It is already extensively applied in tasks with broad real-world applications, e.g., image classification (Frome et al., 2013; Xian et al., 2019), semantic segmentation (Bucher et al., 2019; Ding et al., 2022), video understanding (Xu et al., 2021; Yang et al., 2022), 3D generation (Jain et al., 2022; Xu et al., 2023), etc., which also contributes significantly to the robust development of Large Language Models (LLMs) (Wei et al., 2021; Kojima et al., 2022) and Embodied AI (Varley et al., 2024; Huang et al., 2022).
@@ 3.2. Cross-modal Feature Extraction
3.3. Vision Instance-level Contrastive Learning
@@
where $\tilde{v}$, $\tilde{v}^{+}$ and $\tilde{v}^{-}$ represent a candidate positive sample, its positive sample and negative sample respectively. And we adopt a strategy that adjusts for intra-category visual variability through a similarity-based selection of positive samples. Given a batch, the similarity score $D(\tilde{v},\tilde{v}^{+})$ is computed. If no intra-category sample resemble the candidate, we then identify the similar samples across the batch to serve as the positive samples, irrespective of category, based on the similarity score. This approach enables the model to maintain category coherence despite visual discrepancies.
@@ 3.5. Evidential deep learning
(9) $c_{p}(\omega^{A},\omega^{B})=1/2\sum\nolimits_{k=1}^{K}|p_{k}^{A}-p_{k}^{B}|,$
@@
As illustrated in Figure(2), we treat the outputs of VGT and AGT as evidence vectors, which typically involve issues of ambiguous recognition. Employing EDL allows us to precisely quantify these uncertainties, thereby deriving accurate recognition results. For each instance $\{\mathbf{x}_{n}^{m}\}_{m=1}^{M}$, the modality count $M$ encapsulates two modalities in our bidirectional grounding Transformer, namely visual-to-attribute and attribute-to-visual. The network computes Dirichlet distribution parameters $\bm{\alpha}_{n}^{m}=\mathbf{e}_{n}^{m}+\mathbf{1}$, where $\mathbf{e}_{n}^{m}=f_{\theta}^{m}(\mathbf{x}_{n}^{m})$ is the predicted evidence vector, with $f_{\theta}^{m}$ denoting the modality-specific transformation function. The uncertainty mass derived as $u_{n}^{m}=\frac{K}{\sum_{k=1}^{K}(\alpha_{k,n}^{m})}$, where $K=|\mathcal{C}|$. Adapting to unimodal evidence-based classification, the traditional cross-entropy loss is intricately tailored for compatibility with this framework:
@@
$=\sum\nolimits_{j=1}^{K}y_{nj}\left(\psi\left(S_{n}\right)-\psi\left(\alpha_{nj}^{m}\right)\right),$
where $\mathcal{L}_{ACE}(\bm{\alpha}_{n}^{m})$ denotes the unimodal adaptive cross-entropy loss for the parameters $\bm{\alpha}_{n}^{m}$ of the Dirichlet distribution for a single instance $n$. Utilizing the digamma function $\psi$, the integral is simplified to the expectation of the logarithm of predicted probabilities, where $S_{n}$ represents the sum of Dirichlet parameters for instance $n$, reflecting the total evidence across all classes. The objective of this adaptive loss function is to adjust the network’s output parameters to accurately represent the inherent uncertainty in predictions, enabling the network to make confident predictions when evidence is ample and maintain a degree of uncertainty when evidence is scarce.
@@
(12) $\mathcal{L}_{KL}\left(\bm{\alpha}_{n}^{m}\right)$ $=KL\left[D\left(\bm{p}_{n}\mid\tilde{\bm{\alpha}}_{n}^{m}\right)\|D\left(\bm{p}_{n}\mid\mathbf{1}\right)\right]$
@@
where $D\left(p_{n}\mid\mathbf{1}\right)$ represents the uniform Dirichlet distribution, ${\bm{\tilde{\alpha}_{n}}^{m}}=\mathbf{y_{n}}+(\mathbf{1}-\mathbf{y_{n}})\odot\bm{\alpha_{n}^{m}}$ denotes the Dirichlet parameters after excluding non-misleading evidence from the predicted parameters $\bm{\alpha}_{n}^{m}$ for the $n$-th instance, and $\Gamma(\cdot)$ signifies the gamma function.
@@
(13) $\mathcal{L}_{ACC}\left(\bm{\alpha}_{n}^{m}\right)=L_{ACE}\left(\bm{\alpha}_{n}^{m}\right)+\lambda_{t}L_{KL}\left(\bm{\alpha}_{n}^{m}\right),$
@@ 4. Experiments
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ 4.2. Ablation Studies
In the ablation study depicted in the first image, the effectiveness of various components of the CREST model is evaluated on CUB and SUN datasets. The study illustrates the importance of each component to the model’s performance in both GZSL and CZSL.
@@ 4.3. Hyperparameter Analysis
The parameter tuning for the CREST model indicates a clear optimum range for both $\lambda_{CAL}$ and $\lambda_{EDL}$. Performance peaks at moderate values of $\lambda_{CAL}$ before declining, signifying its critical role in balancing GZSL and CZSL outcomes. The influence of $\lambda_{EDL}$ appears more stable, with only a slight drop at high values, suggesting its robust contribution to the model’s consistent performance across diverse visual tasks. These findings highlight CREST’s ability to maintain accuracy while effectively generalizing to new categories, marking its strengths in a zero-shot learning context.

### v4

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28–November 1, 2024; Melbourne, VIC, Australia.booktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28–November 1, 2024, Melbourne, VIC, Australiaisbn: 979-8-4007-0686-8/24/10doi: 10.1145/3664647.3681629ccs: Computing methodologies Learning paradigms22footnotetext: Corresponding author.
@@ 1. Introduction
Humans frequently possess the talent to grasp novel concepts relying on prior experience without the need to see them beforehand. For instance, a peacock is commonly known as a bird with a colorful fan-shaped tail; if individuals have previous knowledge of birds and fans, they can quickly identify a peacock. However, unlike humans, widely used and studied supervised deep learning models are typically limited to classifying samples belonging to categories seen during training, lacking the capacity to handle samples from unseen categories during training, thus lacking generality and flexibility. Therefore, to further advance Artificial General Intelligence (AGI) (Bostrom, 2020) and achieve true implementation, Zero-Shot Learning (ZSL) was introduced to identify new classes by leveraging inherent semantic relationships during learning (Larochelle et al., 2008; Palatucci et al., 2009; Lampert et al., 2009, 2013; Fu et al., 2017). It is already extensively applied in tasks with broad real-world applications, e.g., image classification (Frome et al., 2013; Xian et al., 2019), semantic segmentation (Bucher et al., 2019; Ding et al., 2022), video understanding (Xu et al., 2021; Yang et al., 2022; Chen et al., 2024a), 3D generation (Jain et al., 2022; Xu et al., 2023; Chen et al., 2024b), etc., which also contributes significantly to the robust development of Large Language Models (LLMs) (Wei et al., 2021; Kojima et al., 2022) and Embodied AI (Varley et al., 2024; Huang et al., 2022).
@@ 3.2. Cross-modal Feature Extraction
3.3. Visual Instance-level Contrastive Learning
@@
where $s$, $\tilde{v}$, $\tilde{v}^{+}$ and $\tilde{v}^{-}$ represent the input sentences from language side, a candidate positive sample, its positive sample and negative sample respectively. And we adopt a strategy that adjusts for intra-category visual variability through a similarity-based selection of positive samples. Given a batch, the similarity score $D(\tilde{v},\tilde{v}^{+})$ is computed. If no intra-category sample resemble the candidate, we then identify the similar samples across the batch to serve as the positive samples, irrespective of category, based on the similarity score. This approach enables the model to maintain category coherence despite visual discrepancies.
@@ 3.5. Evidential deep learning
(9) $c_{p}(\omega^{A},\omega^{B})=\frac{\sum\nolimits_{k=1}^{K}|p_{k}^{A}-p_{k}^{B}|}{2},$
@@
As illustrated in Figure(2), we treat the outputs of VGT and AGT as evidence vectors, which typically involve issues of ambiguous recognition. Employing EDL allows us to precisely quantify these uncertainties, thereby deriving accurate recognition results. For each instance $\{\mathbf{x}_{n}^{m}\}_{m=1}^{M}$, the modality count $M$ encapsulates two modalities in our bidirectional grounding Transformer, namely visual-to-attribute and attribute-to-visual. The network computes Dirichlet distribution parameters $\bm{\alpha_{n}^{m}}=\bm{{e}_{n}^{m}}+\mathbf{1}$, where $\mathbf{e}_{n}^{m}=f_{\theta}^{m}(\mathbf{x}_{n}^{m})$ is the predicted evidence vector, with $f_{\theta}^{m}$ denoting the modality-specific transformation function. The uncertainty mass derived as $u_{n}^{m}=\frac{K}{\sum_{k=1}^{K}(\alpha_{k,n}^{m})}$, where $K=|\mathcal{C}|$. Adapting to unimodal evidence-based classification, the traditional cross-entropy loss is intricately tailored for compatibility with this framework:
@@
$=\sum\nolimits_{j=1}^{K}y_{nj}\left(\psi\left(S_{n}^{m}\right)-\psi\left(\alpha_{nj}^{m}\right)\right),$
where $\mathcal{L}_{ACE}(\bm{\alpha}_{n}^{m})$ denotes the unimodal adaptive cross-entropy loss for the parameters $\bm{\alpha}_{n}^{m}$ of the Dirichlet distribution for a single instance $n$. Utilizing the digamma function $\psi$, the integral is simplified to the expectation of the logarithm of predicted probabilities, where $S_{n}^{m}$ represents the sum of Dirichlet parameters for instance $n$, reflecting the total evidence across all classes. The objective of this adaptive loss function is to adjust the network’s output parameters to accurately represent the inherent uncertainty in predictions, enabling the network to make confident predictions when evidence is ample and maintain a degree of uncertainty when evidence is scarce.
@@
(12) $\mathcal{L}_{KL}\left(\bm{\alpha_{n}^{m}}\right)$ $=KL\left[D\left(\bm{p_{n}^{m}}\mid{\bm{\tilde{\alpha}_{n}^{m}}}\right)\|D\left(\bm{p}_{n}^{m}\mid\mathbf{1}\right)\right]$
@@
where $D\left(\bm{p_{n}^{m}}\mid\mathbf{1}\right)$ represents the uniform Dirichlet distribution, ${\bm{\tilde{\alpha}_{n}^{m}}}=\mathbf{y_{n}}+(\mathbf{1}-\mathbf{y_{n}})\odot\bm{\alpha_{n}^{m}}$ denotes the Dirichlet parameters after excluding non-misleading evidence from the predicted parameters $\bm{\alpha_{n}^{m}}$ for the $n$-th instance, and $\Gamma(\cdot)$ signifies the gamma function.
@@
(13) $\mathcal{L}_{ACC}\left(\bm{\alpha_{n}^{m}}\right)=\mathcal{L}_{ACE}\left(\bm{\alpha_{n}^{m}}\right)+\lambda_{t}\mathcal{L}_{KL}\left(\bm{\alpha_{n}^{m}}\right),$
@@ 4.2. Ablation Studies
In the ablation study depicted in Table 4.2, the effectiveness of various components of the CREST model is evaluated on CUB and SUN datasets. The study illustrates the importance of each component to the model’s performance in both GZSL and CZSL.
@@ 4.3. Hyperparameter Analysis
The parameter tuning for the CREST model indicates a clear optimum range for both $\lambda_{CAL}$ and $\lambda_{EDL}$ in Figure 8. Performance peaks at moderate values of $\lambda_{CAL}$ before declining, signifying its critical role in balancing GZSL and CZSL outcomes. The influence of $\lambda_{EDL}$ appears more stable, with only a slight drop at high values, suggesting its robust contribution to the model’s consistent performance across diverse visual tasks. These findings highlight CREST’s ability to maintain accuracy while effectively generalizing to new categories, marking its strengths in a zero-shot learning context.